### 27 Mar 2026 v1

# Complexity of Quadratic Bosonic Hamiltonian Simulation:

Abstract
The computational complexity of simulating the dynamics of physical quantum systems is a central question at the interface of quantum physics and computer science. In this work, we address this question for the simulation of exponentially large bosonic Hamiltonians with quadratic interactions. We present two results: First, we introduce a broad class of quadratic bosonic problems for which we prove that they are $\mathsf{BQP}$-complete. Importantly, this class includes two known $\mathsf{BQP}$-complete problems as special cases: Classical oscillator networks and continuous-time quantum walks. Second, we show that extending the aforementioned class to even more general quadratic Hamiltonians results in a $\mathsf{PostBQP}$-hard problem. This reveals a sharp transition in the complexity of simulating large quantum systems on a quantum computer, as well as in the difference in complexity between their simulation on classical and quantum computers.
Quadratic bosonic (i.e., Gaussian) Hamiltonians sit between quantum and classical systems. On the one hand, they exhibit clear quantum features, such as sub-classical noise due to squeezing. On the other hand, the resulting Heisenberg equations are closed on the level of single bosonic operators, and thus only couple moments of the same order. Therefore, given moments of some order RR—which can be encoded in a vector of dimension $d=O(M^{R})$—the dynamics of a system of $M$ bosonic modes can be simulated efficiently (i.e., with resources ${\mathsf{poly}}(M)$). In particular, this implies that these systems cannot capture the full power of quantum computing (unless $\mathsf{BQP}$$\mathsf{BPP}$, or moments of unbounded order are involved).
That the dynamics of these systems is captured by a linear first-order differential equation of size $d=O(M^{R})$ raises an intriguing question: Can we gain an advantage by simulating these restricted quantum systems of a fully fledged quantum computer, given that the problem can be represented using $\log d=O(R\log M)$ qubits? For the related problem of quadratic fermionic Hamiltonians, this has indeed been demonstrated [1]. As for bosons, Babbush et al. [2] have recently shown how to simulate exponentially large classical systems with quadratic Hamiltonians on a quantum computer, and that this problem is, in fact, $\mathsf{BQP}$-complete. As the evolution equations for the first moments are identical for the classical and quantum Hamiltonian, this implies $\mathsf{BQP}$-completeness also for the corresponding class of bosonic quantum Hamiltonians. A seemingly unrelated problem which allows for exponential compression and is $\mathsf{BQP}$-complete are continuous-time quantum walks, which have long been used as a way to reason about quantum computing [3, 4]. Remarkably, quantum walks can also be understood as a special instance of bosonic Gaussian Hamiltonians—namely, a quadratic hopping Hamiltonian acting on the subspace with one boson.
That for these two simple and distinct classes of Gaussian Hamiltonians, exponential-size instances are $\mathsf{BQP}$-complete, raises questions about the computational complexity of general bosonic Gaussian Hamiltonians: What is the most general class of bosonic Gaussian Hamiltonians for which exponentially large instances are $\mathsf{BQP}$-complete—ideally including both of the aforementioned problems as special cases? And are there limits to this—are all bosonic Gaussian Hamiltonians exponentially compressible, or are there classes of quadratic bosonic Hamiltonians which—despite the fact that they are classically efficiently simulatable—no longer allow for an exponential advantage when simulated on a quantum computer, and for which the exponential gap between classical and quantum simulation therefore breaks down?
In this work, we introduce a class of quadratic bosonic Hamiltonians without position-momentum coupling, which we term inertially coupled bosons, and prove that simulating their dynamics of exponentially many modes is $\mathsf{BQP}$-complete. This class on the one hand contains systems of quantum harmonic oscillators and thus subsumes the classical oscillator networks studied in Ref. [2]. At the same time, we show that the Hamiltonians governing continuous-time quantum walks can be realized as a particular form of inertial coupling, evolving a state with a single bosonic excitation. Inertially coupled bosons thus provide a single framework which unifies and generalizes classical oscillator dynamics and quantum walk computation within the same complexity theoretic setting. We further extend the class of inertially coupled bosons to all particle number–conserving quadratic (“hopping”) Hamiltonians [^1] In contrast to quantum walks and inertially coupled bosons, these Hamiltonians may include position–momentum couplings. while retaining $\mathsf{BQP}$-completeness.
Subsequently, we proceed to address the question whether all quadratic Hamiltonians can be placed inside $\mathsf{BQP}$, and answer it in the negative: We show that allowing additional quadratic terms beyond inertial couplings and hopping interactions (the classes discussed above) leads to a sudden jump in the computational complexity of the problem. Specifically, the inclusion of such additional terms makes the problem $\mathsf{PostBQP}$-hard (equivalently, $\mathsf{PP}$-hard), which, remarkably, is identical to the complexity of classically simulating the very same system. This establishes a sharp boundary between quantumly simulable and computationally hard regimes within quadratic bosonic systems, and at the same time reveals a surprising jump in the complexity gap between the classical and quantum simulation of quadratic bosonic Hamiltonians.
[…]
Figure 1: Venn diagram depicting inclusion of different classes of Hamiltonians in the class of inertially coupled bosons, cf. Eq. (5). We define the class of inertially coupled bosons as a subset of quadratic bosonic Hamiltonians and prove that their simulation can be formulated as a $\mathsf{BQP}$-complete decision problem. This unifies previous proofs for $\mathsf{BQP}$-completeness of classical and quantum oscillators as well as quantum walks and general hopping Hamiltonians extending the class of inertially coupled bosons. Shades of blue indicate $\mathsf{BQP}$-completeness and Hamiltonians in red define a $\mathsf{PostBQP}$-hard problem, in general.
Inertially Coupled Bosons.—We consider a system of $M=2^{m}$ bosonic modes with position and momentum operators $q_{j}$ and $p_{j}$ (with mode label $j$) satisfying the bosonic commutation relations $\left[q_{j},p_{k}\right]=i\delta_{jk}\mathds{1}$, $\left[q_{j},q_{k}\right]=\left[p_{j},p_{k}\right]=0$. A system of inertially coupled bosons is defined by the quadratic Hamiltonian
$H=\frac{1}{2}\sum_{j,k}(C_{jk}p_{j}p_{k}+A_{jk}q_{j}q_{k}),$ (1)
where $A,C\in\operatorname{Mat}(M\times M,\mathds{R})$ are symmetric $d$-sparse matrices, i.e. each row and column contains at most $d$ nonzero entries. For now, we require $A$ and $C$ to be positive semidefinite (PSD), amounting to a lower semibounded spectrum of the bosonic Hamiltonian. Further details, including extensions beyond the PSD case, are provided in Appendix C.2.
For our construction, we impose a sign convention on the elements of $A$, motivated by a potential energy arising from spring-like coupling, i.e., of the form
$\frac{1}{2}\sum_{j,k}k_{jk}(q_{j}-q_{k})^{2}+\frac{1}{2}\sum_{j}k_{jj}q_{j}^{2}$ (2)
with positive coupling constants $k_{jk}=-A_{jk}$ for interactions between the $j^{\text{th}}$ and $k^{\text{th}}$ boson, and positive coefficients
$k_{jj}=A_{jj}+\sum_{j^{\prime}\neq j}A_{jj^{\prime}}$ (3)
for the coupling to a reference point, [^2] Nonzero $k_{jj}$ restrict the motion of the bosons and prevent arbitrary displacements from the reference point. This can model weak boundaries of an effective confining volume. and analogously for $C$ as arising from a kinetic energy of the form
$\frac{1}{2}\sum_{j,k}l_{jk}(p_{j}-p_{k})^{2}+\frac{1}{2}\sum_{j}l_{jj}p_{j}^{2},$ (4)
with coupling constants $l_{jk}\geq 0$ for all $j,k$. This is equivalent to requiring that $A$ and $C$ are Laplacian-type stiffness matrices, which are positive semidefinite and diagonally dominant. Note that this specific sign convention (beyond $A$ and $C$ being PSD) is not required for $\mathsf{BQP}$-completeness, but simplifies the construction of the simulation algorithm.
With this sign convention, the model corresponds to the quantum analog of a classical Hamiltonian derived from couplings via positive spring constants, now acting in both position and momentum space. Special cases include a quantum analog of a spring network when $C=M^{-1}$ is the inverse of the (diagonal) mass matrix $M$, and (possibly multiparticle) quantum walks when $A=C$, which we discuss in detail later.
Off-diagonal entries of $C$, corresponding to coupling between the momenta of different bosons, can be interpreted as inertial coupling. Such couplings naturally arise in constrained classical systems, for example in the classical double pendulum.
Note that the Hamiltonian in Eq. (1) is in general not particle-number conserving. Yet, it remains 2-local (i.e. each term has support on at most two modes). We will later show that the $\mathsf{BQP}$-complete class can be extended to include general particle-number–preserving quadratic Hamiltonians, including those with complex hopping amplitudes. For clarity of presentation, we focus here on the inertially coupled case corresponding to Eq. (1).
Going beyond inertially coupled bosons, the most general quadratic bosonic Hamiltonian,
$H=\frac{1}{2}\begin{pmatrix}q&p\end{pmatrix}\begin{pmatrix}A&E\\ E^{T}&C\end{pmatrix}\begin{pmatrix}q\\ p\end{pmatrix}\$, (5)
includes mixed position–momentum couplings described by a matrix $E_{jk}=E^{+}_{jk}+iE^{-}_{jk}$, which can be split into a real symmetric component $E^{+}$ and a real antisymmetric component $iE^{-}$. These mixed $q$-$p$ terms include both general hopping interactions and squeezing-type terms. As we show later, these terms enable computational power beyond $\mathsf{BQP}$.
Simulation of inertially coupled bosons.—The problem that we consider is to simulate time evolution for a bosonic system of size $M=2^{n}$ bosons with sufficiently sparse inertial coupling for a time $t={\mathsf{poly}}(n)$. The task is to estimate sums of squared expectation values $\sum_{O}\left\lvert\left\langle O(t)\right\rangle\right\rvert^{2}$ of time-evolved operators, with $O$ from a given class of $R$-local (i.e., $R$-body) observables, where at most ${\mathsf{poly}}(n)$ of the initial values $\left\langle q_{j}(0)\right\rangle$ and $\left\langle p_{j}(0)\right\rangle$, as well as of the higher moments up to $R^{\text{th}}$ order, are non-zero.
We prove that this problem is $\mathsf{BQP}$-complete. To this end, we follow an encoding scheme that has been used for classical oscillators [2] and more recently for shadow Hamiltonian simulation [5]. Encoding expectation values of a suitable set of observables $O_{a}$ as a quantum state via $\left|\psi\right\rangle=\sum_{a}\left\langle O_{a}\right\rangle\left|a\right\rangle$ allows for a simulation in the Heisenberg picture by integrating the von Neumann equation. Although the resulting effective time propagator for $\left|\psi\right\rangle$ does not need to be unitary, through the right choice of the operators $O_{a}$ defining the encoding one can achieve unitary dynamics. We show how to obtain such a Hamiltonian formulation for the class of inertially coupled bosonic systems by choosing $O_{a}$ as $R$-local correlation functions of the canonical operators $q_{j}$ and $p_{j}$, whose dynamics are closed under quadratic Hamiltonians. This allows for an exponential compression of the data for $M=2^{m}$ bosonic modes into $O(m)$ qubits.
We denote all canonical operators by $z=(q_{1},...,q_{M},p_{1},...,p_{M})^{T}$. We index the canonical operators collectively by $a$, and distinguish position and momentum operators by $j$ and $\bar{j}$, i.e. $z_{j}=q_{j}$ and $z_{\bar{j}}=p_{j}$. The time evolution of such a system is determined by the von Neumann equation, which can be written in the Heisenberg picture as
$\partial_{t}z=i\left[H,z\right]=\begin{pmatrix}0&C\\ -A&0\end{pmatrix}z.$ (6)
[…]
$B=$ $\sum_{j<k}\sqrt{-A_{jk}}(\left|j\right\rangle-\left|k\right\rangle)\left\langle j,k\right|$
$+\sum_{j}\sqrt{A_{jj}+\sum_{k\neq j}A_{jk}}\left|j\right\rangle\left\langle j,j\right|,$ (7)
and $D$ analogously with $A\leftrightarrow C$. We introduce new coordinates via $z_{\kappa}=i\sum_{j}B^{\dagger}_{\kappa j}q_{j}$ and $z_{\bar{\kappa}}=\sum_{j}D^{\dagger}_{\kappa j}p_{j}$, which we will distinguish from the original mode indices $z_{k},z_{\bar{k}}$ by greek subscripts. Alternatively, we use $\alpha\in~\left\{1,...,\binom{M+1}{2},\bar{1},...,\overline{\binom{M+1}{2}}\right\}$ to denote both $(j,k)$ and $(\bar{j},\bar{k})$ with $j\leq k$. This redundant encoding only leads to constant factors in memory, while allowing for a reconstruction of the original degrees of freedom $z_{a}$.
Now consider the following decision problem.
Problem 1 (Simulation of Inertially Coupled Bosons).
An instance of the problem consists of:
(i) Oracle access to a $d$-sparse Hamiltonian $H$ of the form Eq. (1) acting on $M$ bosonic modes (where we can query the non-zero entries of any row of $H$)
(ii) A list of $\mathcal{O}({\mathsf{polylog}}(M))$ nonzero initial moments
$Z_{\alpha_{1}\ldots\alpha_{r}}:=\langle z_{\alpha_{1}}(0)\cdots z_{\alpha_{r}}(0)\rangle,\quad 1\leq r\leq R,$ (8)
with $R=\mathcal{O}(1)$, each specified to polynomial precision
(iii) A population $\zeta(I_{0})$ over an efficiently characterizable subset $I_{0}$:
$\zeta(I_{0};t)=\sum_{r=1}^{R}\sum_{(\alpha_{1},\ldots,\alpha_{r})\in I_{0}}\left\lvert\left\langle\prod_{i=1}^{r}z_{\alpha_{i}}(t)\right\rangle\right\rvert^{2},$ (9)
(iv) A final time $t_{f}=\mathcal{O}({\mathsf{polylog}}(M))$ and real numbers $a>b$ satisfying $a-b=\Omega(1/{\mathsf{polylog}}(M))$.
The operators evolve according to the Heisenberg equations of motion generated by $H$. The task is to decide whether $\zeta(I_{0};t_{f})>a$ or $\zeta(I_{0};t_{f})<b$, promised that one of the two cases holds.
The restriction on the number of nonzero moments ensures that the input admits a polynomial-size classical description, and is equivalent to assuming access to a polynomial-size quantum circuit preparing the relevant initial state.
Our first result is as follows:
Theorem 1.
Problem 1 is $\mathsf{BQP}$-complete.
$\mathsf{BQP}$-hardness follows immediately since the classical oscillator dynamics of [2] arises as the $R=1$ sector of Problem 1. The equations of motion for the first moments $\left\langle q_{j}\right\rangle$ and $\left\langle p_{j}\right\rangle$ (corresponding to $R=1$) of a quadratic quantum Hamiltonian are identical to those of the corresponding classical Hamiltonian (1). For a diagonal mass matrix $C$, this describes a system of masses coupled by springs, which has been shown to be $\mathsf{BQP}$-hard [2].
To prove containment in $\mathsf{BQP}$, we proceed in three steps: (i) Since $R$-local observables are closed under the dynamics governed by quadratic Hamiltonians, we first encode the initial values, $\left\langle z_{\alpha_{1}}(0)\cdots z_{\alpha_{r}}(0)\right\rangle$, into a quantum state $\left|\psi\right\rangle$, which is efficiently preparable given a polynomial description of the initial moments. (ii) We construct an effective Hamiltonian $H_{0}$ which governs the time evolution of $\left|\psi\right\rangle$. It is sparse and can be efficiently simulated on a quantum computer, using for instance a block-encoding of $H_{0}$. (iii) The readout of the form of Eq. (9) can be estimated by measurements in the computational basis.
(i) Encoding.
In order to simulate quantum oscillators in the Heisenberg picture, we encode the expectation values of all products of up to $R$ generalized coordinates $z_{\alpha}$ into a quantum state [^3] The information about the same generalized coordinates $z_{a}$ is stored in multiple basis states $\left|\alpha_{1}...\alpha_{r}\right\rangle$ by redundancy in $B$ and $D$. There are ways to reduce memory by constant factors, but we stick to this encoding for the sake of clarity.
$\left|\psi(t)\right\rangle=\frac{1}{N}\sum_{r=1}^{R}\sum_{\alpha_{1},...,\alpha_{r}=1}^{2\binom{M+1}{2}}\left\langle\prod_{i=1}^{r}z_{\alpha_{i}}(t)\right\rangle\left|\alpha_{1}...\alpha_{r}\right\rangle\ ,$ (10)
which contains all expectation values of $R$-local monomials in the generalized coordinates at a given time $t$ (which will allow for the simulation of their dynamics). [^4] We can directly read off the coordinate $z_{a}$ with $a=j$ ($a=\bar{j}$) from the amplitude $\left\langle z_{(a,a)}\right\rangle$, given that its interaction to the wall given by $k_{jj}$ ($l_{jj}$) is nonzero. Otherwise, the measurement of the relative coordinates $z_{j,k}$ ($z_{\bar{j},\bar{k}}$) is necessary to access $z_{a}$, see Appendix A. $\left|\psi(t)\right\rangle$ can be encoded using $\log\left(\sum_{r=1}^{R}\left(2\binom{M+1}{2}\right)^{r}\right)=\mathcal{O}(R\log(M))$ qubits using amplitude encoding. Such a state can be efficiently prepared for polynomially many nonzero amplitudes [6, 7], such as for the initial state in Problem 1.
(ii) Hamiltonian Encoding of Dynamics.
The dynamics of the amplitudes $\left\langle\prod_{i=1}^{r}z_{\alpha_{i}}(t)\right\rangle$ are governed by the von Neumann equation (6), which, using the Leibniz rule for derivations, straightforwardly evaluates to
$\partial_{t}\left\langle z_{\alpha_{1}}...z_{\alpha_{r}}\right\rangle$ $=\sum_{j=1}^{r}\left\langle z_{\alpha_{1}}\cdots z_{\alpha_{j-1}}(iH_{0}z)_{\alpha_{j}}z_{\alpha_{j+1}}\cdots z_{\alpha_{r}}\right\rangle$
$\text{with}\quad H_{0}=\begin{pmatrix}0&B^{\dagger}D\\ D^{\dagger}B&0\end{pmatrix}$. (11)
Written more compactly, this is an equation for the rank-$r$ tensor
$\partial_{t}\left\langle z^{\otimes r}\right\rangle=i\sum_{j}H_{0}^{(j)}\left\langle z^{\otimes r}\right\rangle,$ (12)
using the shorthand $H_{0}^{(j)}$, which has support only on the $j^{\text{th}}$ tensor factor. Since for each $r=1,\dots,R$, Eq. (12) describes a Hamiltonian evolution of the corresponding sector of $\left|\psi(t)\right\rangle$, Eq. (10), the encoded state $\left|\psi(t)\right\rangle$ evolves according to a Schrödinger equation.
Starting from an oracle for the $d$-sparse coupling matrices $A$ and $C$, one can build $d$-sparse oracles for $B,B^{\dagger}$ and $D,D^{\dagger}$ [8]. With this, it is straightforward to construct an oracle for the $rd$-sparse Hamiltonian on the right hand side of Eq. (12). Evolution under sparse Hamiltonians can be efficiently simulated, for instance using one of the methods in [9, 10, 11, 12]. In direct analogy to [2], we can ensure a final error $\varepsilon$ using $Q=\mathcal{O}(tKd\left\lVert H_{0}\right\rVert_{\rm max}+\log(1/\varepsilon))$ queries of $H_{0}$ and a number of gates $G=\mathcal{O}(Q\log^{2}(\frac{MQ}{\varepsilon}))$.
(iii) Measurement.
The read-out of the problem is the estimation of the sum of the squared amplitudes
$\left\lvert\left\langle\prod_{i=1}^{r}z_{\alpha_{i}}\right\rangle\right\rvert^{2}=\left\lvert\left\langle\alpha_{1}...\alpha_{r}|\psi(t)\right\rangle\right\rvert^{2}N^{2}$ (13)
over $(\alpha_{1},...,\alpha_{r})\in I_{0}$ for an efficiently characterizable index set $I_{0}$. The normalization factor $N$ from Eq. (10) is set by the initial conditions and remains unchanged under unitary time evolution.
We use an ancilla qubit to encode those basis states that belong to $I_{0}$ via
$\left|\alpha_{1},...,\alpha_{r}\right\rangle\left|0\right\rangle\mapsto\left|\alpha_{1},...,\alpha_{r}\right\rangle\left|f(\alpha_{1},...,\alpha_{r})\right\rangle$ (14)
with $f(\alpha_{1},...,\alpha_{r})=1$ if $(\alpha_{1},...,\alpha_{r})\in I_{0}$ and 0 else. Finally, the probability to measure the ancilla qubit in the state $\left|1\right\rangle$ yields the desired estimate, solving problem 1. It is also possible to measure the reconstructed amplitudes $\left\langle z_{a}\right\rangle$ including their sign. We discuss this in appendix A.
Quantum Walks are Inertially Coupled Bosons.—We now relate inertially coupled bosonic systems to continuous-time quantum walks, a canonical model of quantum computation known to be $\mathsf{BQP}$-complete [3, 4]. It was shown in [8] that classical oscillator networks can directly simulate such quantum walks, thereby recovering the $\mathsf{BQP}$-hardness result of [2]. Both descriptions naturally fall within the class of Hamiltonians for inertially coupled bosons.
[…]
$H$ $=\frac{1}{2}\sum_{j,k}\tilde{T}_{jk}\left(q_{j}-ip_{j}\right)\left(q_{k}+ip_{k}\right)$
$=\frac{1}{2}\sum_{j,k}\tilde{T}_{jk}\left(q_{j}q_{k}+p_{j}p_{k}+i(q_{j}p_{k}-p_{j}q_{k})\right)\>.$ (15)
[…]
$H$ $=\frac{1}{2}\sum_{j,k}\tilde{T}_{jk}\left(q_{j}q_{k}+p_{j}p_{k}\right)\ .$ (16)
This is precisely the Hamiltonian of inertially coupled bosons, Eq. (1), with positive semidefinite coupling matrices $A=C=\tilde{T}$.
Note that $\tilde{T}$, and thus $A$ and $C$, have precisely the structure required for a network of harmonically coupled oscillators: Its off-diagonal elements are non-positive,
$A_{jk}=C_{jk}=-k_{jk}\leq 0\quad(j\neq k)\ ,$ (17)
while the diagonal entries satisfy
$A_{jj}=C_{jj}=k_{jj}+\sum_{k\neq j}k_{jk}\ ,$ (18)
with $k_{jk}\geq 0$ interpreted as spring constants between modes $j$ and $k$ making it a Laplacian-type stiffness matrix.
More generally, number-preserving hopping Hamiltonians with complex amplitudes yield additional antisymmetric terms of the form $q_{j}p_{k}-q_{k}p_{j}$ in the $(q,p)$ representation. A detailed characterization of such terms is given in Appendix C.
Simulating general quadratic Hamiltonians is $\mathsf{PostBQP}$-hard.—We now show that extending the bosonic Hamiltonian beyond the $\mathsf{BQP}$-complete subclass identified above fundamentally increases the computational power of the dynamics, elevating the simulation problem to become $\mathsf{PostBQP}$-hard (that is, quantum circuits augmented by the possibility to project onto specific measurement outcomes). The key distinction is that more general quadratic Hamiltonians, beyond those considered up to now, induce a linear evolution on the first moments which is no longer norm-preserving. Such terms enable the exponential amplification or suppression of selected components, which in turn can be used to emulate quantum circuits with postselection.
Theorem 2.
The variant of problem 1 in which $H=\frac{1}{2}\sum_{j,k}H_{jk}z_{j}z_{k}$ is an arbitrary, sparse and efficiently queriable quadratic Hamiltonian is $\mathsf{PostBQP}$-hard.
The proof, given in Appendix B, is based on the Feynman–Kitaev construction, as also used to establish $\mathsf{BQP}$-completeness for classical oscillators [2]. However, by additionally introducing squeezing terms of the form $q_{j}p_{j}$ via $E^{+}$ in Eq. (5), we are able to implement an exponential suppression of the computational branch corresponding to an output qubit in state $\left|0_{P}\right\rangle$ relative to $\left|1_{P}\right\rangle$, thereby approximately realizing the desired postselection on the state $\left|1_{P}\right\rangle$ at the end of the circuit.
Other families of Hamiltonians with $E^{+}=0$ can also implement postselection. As we show in Appendix B.7, negative eigenvalues in $A$ or mixed terms from $E^{-}\neq 0$ can be used to create the same dynamics as in the aforementioned family with $E^{+}\neq 0$.
Conclusions.—We have studied the complexity of simulating the time evolution of quadratic bosonic Hamiltonians, acting on an exponentially large number of modes $M=2^{n}$, when restricted to moments of bounded order $R$. We have identified two regimes: On the one side, we have introduced the class of inertially coupled bosons, that is, systems without coupling between positions and momenta, and shown that simulating its dynamics is $\mathsf{BQP}$}-complete, and that this class encompasses the known $\mathsf{BQP}$-complete problems of simulating quantum walks and classical harmonic oscillators as special cases. We have also shown how to generalize this class such as to encompass all hopping Hamiltonians while retaining $\mathsf{BQP}$-completeness. On the other side, we found that any further generalization beyond this class, but still within the family of bosonic quadratic Hamiltonians, gave rise to a sudden jump in complexity, making the problem $\mathsf{PostBQP}$-hard. Interestingly, the corresponding classical complexity class $\mathsf{PP}$, which equals $\mathsf{PostBQP}$, naturally captures the complexity of simulating arbitrary exponentially large bosonic quadratic Hamiltonians, regardless of the family; it would be interesting to investigate this apparent vanishing of the difference in complexity of classical and quantum simulation more closely.
A noteworthy point is the role played by squeezing in this complexity transition. On the one hand, we find that all hopping Hamiltonians (that is, those without squeezing terms) are inside $\mathsf{BQP}$. On the other hand, the presence of a squeezing term does not immediately imply $\mathsf{PostBQP}$-hardness, as can be seen by e.g. setting $A\neq C$ in Eq. (1). The reason for this can be intuitively grasped already for a single mode: The presence of a squeezing term in the Hamiltonian does not necessarily imply that squeezing builds up over time, as rotating terms can lead to an averaging out of the squeezing. Let us remark that a similar effect of squeezing on complexity has been observed in the context of simulating photonic circuits with particle number-conserving gates: As long as only a bounded number of squeezing operations is allowed, the problem is $\mathsf{BQP}$-complete [13], while it becomes $\mathsf{PostBQP}$-hard once a linear amount of squeezing is permitted [14].
Acknowledgments.—We acknowledge support from the Austrian Science Fund FWF (Grant No. 10.55776/COE1 and 10.55776/F71), by the European Union – NextGenerationEU, and by the European Union’s Horizon 2020 research and innovation programme through Grant No. 863476 (ERC-CoG SEQUAM).
Appendix A Post-Processing of Measurement Data
In this appendix, we extend the discussion on the read-out of quadrature expectation values and discuss how to reconstruct the expectation values $\left\langle z_{a_{1}}\cdots z_{a_{r}}\right\rangle$ from the encoded amplitdues $\left\langle z_{\alpha_{1}}\cdots z_{\alpha_{r}}\right\rangle$ and how to improve efficiency including all samples in computational basis measurements using the example of linear and quadratic expectation values.
Because position coordinates are mixed by $B$ and momentum coordinates by $D$, only specific linear combinations of the generalized coordinates $z_{a}$ are directly accessible through computational-basis measurement. However, the encoding via the indices $\alpha$ is redundant, and expectation values of individual generalized coordinates can be reconstructed from diagonal components. In particular, the coordinates
$\left\langle z_{(j,j)}\right\rangle$ $=\sum_{k=1}^{M}B^{\dagger}_{(j,j)k}\left\langle q_{k}\right\rangle=\sqrt{A_{jj}+\sum_{k\neq j}A_{jk}}\,\left\langle q_{j}\right\rangle$ (19)
directly encode the position of mode $j$, provided that $A_{jj}+\sum_{k\neq j}A_{jk}\neq 0$. Analogous statements hold for momentum coordinates using $C$ and $D$. Consequently, the absolute value of any $r$-local correlation $\big|\langle z_{a_{1}}\cdots z_{a_{r}}\rangle\big|$ can be inferred from the probability of measuring the basis state $\left|\alpha_{1}\ldots\alpha_{r}\right\rangle=\left|(a_{1},a_{1})\ldots(a_{r},a_{r})\right\rangle$, assuming nonvanishing coupling to the reference.
To recover sign information, one can use a superposition of the vacuum component $\left|\mathds{1}\right\rangle$ (corresponding to $r=0$) with the basis state $\left|(a_{1},a_{1})\ldots(a_{r},a_{r})\right\rangle$ obtained through a Hadamard transformation. Measuring in the rotated basis yields probabilities proportional to $\left\lvert 1\pm\left\langle z_{(a_{1},a_{1})}\cdots z_{(a_{r},a_{r})}\right\rangle\right\rvert^{2}$, which allows one to infer the sign of $\left\langle z_{a_{1}}\cdots z_{a_{r}}\right\rangle$ whenever it is polynomially separated from zero.
In practice, one can gain efficiency when estimating all relevant amplitudes by recycling measurement outcomes. As measurements may also yield outcomes corresponding to coordinates $z_{(a,b)}$, which encode relative positions and momenta, these additional samples can be incorporated to reduce estimator variance. We restrict the discussion to the cases $r=1$ and $r=2$ and to positional coordinates $q$, governed by the matrix $A$. The analogous construction for momenta follows by replacing $q\to p$, $A\to C$, and $B\to D$.
Linear expectation values.
For $r=1$, we have
$\left\langle z_{(j,j^{\prime})}\right\rangle$ $=\sqrt{-A_{jj^{\prime}}}\left\langle q_{j}-q_{j^{\prime}}\right\rangle$ (20)
$\left\langle z_{(j,j)}\right\rangle$ $=\sqrt{A_{jj}+\sum_{k\neq j}A_{jk}}\,\left\langle q_{j}\right\rangle.$ (21)
The prefactor $\sqrt{A_{jj}+\sum_{k\neq j}A_{jk}}$ can be computed using a single row query of $A$. Measuring the probability of the computational basis state $\left|(j,j)\right\rangle$ therefore allows to compute
$\left\lvert\left\langle q_{j}\right\rangle\right\rvert=\frac{1}{\sqrt{A_{jj}+\sum_{k\neq j}A_{jk}}}\left\lvert\left\langle z_{(j,j)}\right\rangle\right\rvert\$. (22)
The sign of $\left\langle q_{j}\right\rangle$ is not accessible from computational-basis measurements alone but can be recovered by interference, for example via a Hadamard transformation mixing $q_{j}$ with the vacuum component corresponding to $r=0$. A relative sign between two modes can be retrieved mixing $(q_{j},q_{j^{\prime}})\to(q_{j}+q_{j^{\prime}},q_{j}-q_{j^{\prime}})$.
Quadratic correlations.
Certain second moments are directly accessible. For example,
$\left\lvert\left\langle z_{(j,j^{\prime})}^{2}\right\rangle\right\rvert=\left\lvert A_{jj^{\prime}}\right\rvert\left\langle(q_{j}-q_{j^{\prime}})^{2}\right\rangle,$ (23)
which corresponds to the interaction energy associated with the $(j,j^{\prime})$ coupling. More generally, cross-correlations $\left\langle q_{j}q_{j^{\prime}}\right\rangle$ can be reconstructed provided that both modes have nonvanishing coupling to the reference and to each other. In that case one obtains
$\left\langle q_{j}q_{j^{\prime}}\right\rangle$ $=\frac{1}{2}\left(\frac{1}{A_{jj^{\prime}}}\left\langle z_{(j,j^{\prime})}^{2}\right\rangle+\frac{1}{A_{jj}+\sum_{k\neq j}A_{jk}}\left\langle z_{(j,j)}^{2}\right\rangle+\frac{1}{A_{j^{\prime}j^{\prime}}+\sum_{k\neq j^{\prime}}A_{j^{\prime}k}}\left\langle z_{(j^{\prime},j^{\prime})}^{2}\right\rangle\right),$ (24)
or equivalently,
$\left\langle q_{j}q_{j^{\prime}}\right\rangle$ $=\frac{1}{2\sqrt{A_{jj}+\sum_{k\neq j}A_{jk}}\sqrt{A_{j^{\prime}j^{\prime}}+\sum_{k\neq j^{\prime}}A_{j^{\prime}k}}}\left\langle z_{(j,j)}z_{(j^{\prime},j^{\prime})}\right\rangle.$ (25)
Since $\left\langle z_{\alpha}^{2}\right\rangle\geq 0$, we omit absolute values. These expressions yield the energy expectation values $A_{jj^{\prime}}\left\langle q_{j}q_{j^{\prime}}\right\rangle$, provided $A_{jj}+\sum_{k\neq j}A_{jk}\neq 0$ and similarly for $j^{\prime}$. Extensions to higher-order products and to momentum coordinates follow analogously.
Appendix B $\mathsf{PostBQP}$-hardness of generic Gaussian Hamiltonians (Theorem 2)
In this appendix, we prove Theorem 2 by showing that simulating the first-moment dynamics of generic Gaussian Hamiltonians with inertial couplings and mixed terms from $E^{+}$ is $\mathsf{PostBQP}$-hard. The proof proceeds by an explicit reduction from a $\mathsf{PostBQP}$-complete promise problem to the evolution of a quadratic bosonic Hamiltonian. We start by stating the $\mathsf{PostBQP}$-complete bosonic problem in subsection 1, and then proceed to analyze it in subsections 2–5. Subsection 6 contains a technical result required in the analysis, which is standard but provided for completeness. Finally, subsection 7 discusses alternative $\mathsf{PostBQP}$-hard families of bosonic Hamiltonians.
B.1 PostBQP-completeness Construction
$\mathsf{PostBQP}$ (Postselected Bounded-Error Quantum Polynomial Time) is the class of decision problems solvable by a polynomial-time quantum algorithm that is allowed to postselect on the outcome of a measurement [15]. The natural $\mathsf{PostBQP}$-complete problem is postselected circuit: Consider a quantum register consisting of $n$ qubits – a postselection qubit $P$, an output qubit $Q$ and n2n-2 qubits in a register $R$ – and a uniform quantum circuit $U=\prod_{l=1}^{L}U_{l}$, $L={\mathsf{poly}}(n)$, on those $n$ qubits, where the $U_{l}$ are chosen from the computationally universal gateset $\{\mathrm{Hadamard},\mathrm{Toffoli}\}$. Let $\left|\Psi_{1}\right\rangle:=\tfrac{1}{\delta}\langle 1_{P}\rvert U\left|0_{P}0_{Q}0_{R}\right\rangle$ be the normalized state obtained from evolving $\left|0_{P}0_{Q}0_{R}\right\rangle$ with $U$ and postselecting on the $P$ register being $\left|1_{P}\right\rangle$. The task is then to decide whether $\left\lVert\left\langle 1_{Q}|\Psi_{1}\right\rangle\right\rVert^{2}$ is greater than $2/3$ or smaller than $1/3$, promised that one of the two is the case. Importantly, since superpositions of computational basis states are only created by the Hadamard gates in the circuit, we have that $\delta=\Omega(2^{-L})$.
In the following, we provide a reduction from postselected circuit to the simulation of quadratic bosonic Hamiltonians, as defined in Theorem 2 in the main text. The reduction involves three parts: Dynamics, the initial state, and the read-out. We start by stating the individual parts of the reduction, and analyze it in the subsequent subsections.
(i) Dynamics: We encode the quantum circuit $U$ into a Feynman–Kitaev Hamiltonian
$H_{\rm FK}=\sum_{l=1}^{L+1}\left|l\right\rangle\left\langle l+1\right|\otimes U_{l}^{\dagger}+\left|l+1\right\rangle\left\langle l\right|\otimes U_{l}$ (26)
where for technical reasons, we add one extra time step with $U_{L+1}=\mathds{1}$ at the end. $H_{\mathrm{FK}}$ acts on a tensor product of a “clock register” $\left|l\right\rangle$, $l=1,\dots,L+2$, and the $n+1$ qubit register of the circuit. The idea is that time evolution under $H_{\mathrm{FK}}$ propagates an initial state $\left|l=0\right\rangle\otimes\left|\phi\right\rangle$ by simultaneously propagating the time register and applying the corresponding gate $U_{l}$ to the qubit register, such that the state of the system at any time is a superposition of states of the form $\left|l\right\rangle\otimes(U_{l}\cdots U_{1}\left|\phi\right\rangle)$. We encode the Hilbert space of Eq. (26) in terms of $M=(L+2)2^{n}$ bosonic modes, where we label position and momentum operators by $q^{(l)}_{j_{P},j_{Q},j_{R}}$ and $p^{(l)}_{j_{P},j_{Q},j_{R}}$, respectively, with a layer index $l\in\{1,...,L+2\}$, and qubit indices $j_{P},j_{Q}\in\{0,1\}$ and $j_{R}\in\{0,1\}^{\times(n-2)}$. To define the encoding, consider the dynamics of first moments under the Hamiltonian
$H$ $=\frac{1}{2}\begin{pmatrix}q&p\end{pmatrix}\begin{pmatrix}A&E^{+}\\ E^{+}&C\end{pmatrix}\begin{pmatrix}q\\ p\end{pmatrix}\$, (27)
with
$A$ $=4\mathds{1}-\sum_{l}^{L+1}\left(\left|l\right\rangle\left\langle l+1\right|\otimes U_{l}^{\dagger}+\left|l+1\right\rangle\left\langle l\right|\otimes U_{l}\right)\ ,$ (28)
$C$ $=\mathds{1}\ ,$ (29)
$E^{+}$ $=-\beta\left|L+2\right\rangle\left\langle L+2\right|\otimes\left|1_{P}\right\rangle\left\langle 1_{P}\right|.$ (30)
[…]
(ii) Initial state: The initial conditions of the bosonic system to which we reduce Postselected Circuit are given by $\big\langle q_{j}^{(l)}(0)\big\rangle=0$ and $\big\langle p_{j}^{(l)}(0)\big\rangle=\big\langle\dot{q}_{j}^{(l)}(0)\big\rangle=\delta_{l,1}\delta_{j,0}$.
(iii) Read-out: Recall from Eq. (9) that read-outs involve estimating quantities of the form
$\zeta(I;t):=\sum_{r=1}^{R}\sum_{(\alpha_{1},\ldots,\alpha_{r})\in I}\left\lvert\left\langle\prod_{i=1}^{r}z_{\alpha_{i}}(t)\right\rangle\right\rvert^{2}$ (31)
over an efficiently characterizable index set $I$, up to $1/\mathrm{polylog}(M)=1/\mathrm{poly}(n)$ precision. For our reduction, we need to perform two such read-outs after a suitably chosen time $t_{f}=O(L)$. They are given by the index sets
$I_{\mathrm{num}}=\{(l,\alpha)=(\overline{L+1},\bar{1}_{P}\bar{1}_{Q}\bar{j}_{R})\}\mbox{\quad and\quad}I_{\mathrm{den}}=\{(l,\alpha)=(\overline{L+1},\bar{1}_{P}\bar{b}_{Q}\bar{j}_{R})\}\ ,$ (32)
respectively, each for all possible choices of the labels $\bar{b}_{Q}$ and $\bar{j}_{R}$ (recall that the bar denotes momentum labels; that is, the measurement reads out a sum of expectation values of momenta).
Following these two measurements, we accept exactly if $\zeta(I_{\mathrm{num}};t_{f})/\zeta(I_{\mathrm{den}};t_{f})>1/2$. As we will show in the following, for a suitably chosen time $t=O(L)$, this algorithm accepts exactly if the original postselected circuit instance accepts.
B.2 Formal Solution to the Bosonic Problem
The equations of motion defined in Eqs. (27-30) can be rewritten as the following second order linear differential equation in $q$:
$\big\langle\ddot{q}_{j}^{(l)}\big\rangle$ $=\left[-(A-(E^{+})^{2})\left\langle q\right\rangle\right]_{j}^{(l)}=:\left[-\tilde{A}\left\langle q\right\rangle\right]_{j}^{(l)},$ (33)
with $\tilde{A}:=A-(E^{+})^{2}$, which ceases to be positive semidefinite for large enough $\beta$. The momentum coordinates $p_{j}^{(l)}$ can be transformed into velocities $\dot{q}_{j}^{(l)}$ via
$\big\langle\dot{q}_{j}^{(l)}\big\rangle$ $=\big\langle{p_{j}^{(l)}}\big\rangle-\delta_{l,L+2}\delta_{j_{P},1}\beta\big\langle{q_{j}^{(l)}}\big\rangle\ .$ (34)
Correspondingly, the initial state $\big\langle p_{j}^{(l)}(0)\big\rangle=\delta_{l,1}\delta_{j,0}$ translates to $\big\langle\dot{q}_{j}^{(l)}(0)\big\rangle=\delta_{l,1}\delta_{j,0}$, and as before $\big\langle q_{j}^{(l)}(0)\big\rangle\!=\!0$. [^5] Equation (34) is also the reason why we added an extra time step $U_{L+1}=\mathds{1}$, since this way, at time $l=L+1$, $\big\langle\dot{q}_{j}^{(L+1)}\big\rangle=\big\langle{p_{j}^{(L+1)}}\big\rangle$.
The general solution to Eq. (33) reads
$\left\langle\dot{q}(t)\right\rangle+i\sqrt{\tilde{A}}\left\langle q(t)\right\rangle=e^{it\sqrt{\tilde{A}}}\left(\left\langle\dot{q}(0)\right\rangle+i\sqrt{\tilde{A}}\left\langle q(0)\right\rangle\right),$ (35)
with initial conditions $\left\langle\dot{q}(0)\right\rangle$ and $\sqrt{\tilde{A}}\left\langle q(0)\right\rangle$. Using the initial condition $\big\langle q_{j}^{(l)}(0)\big\rangle=0$, this simplifies to
$\left\langle q(t)\right\rangle=\sqrt{\tilde{A}}^{-1}\sin\left(\sqrt{\tilde{A}}t\right)\left\langle\dot{q}(0)\right\rangle,\qquad\text{and}\qquad\left\langle\dot{q}(t)\right\rangle=\cos\left(\sqrt{\tilde{A}}t\right)\left\langle\dot{q}(0)\right\rangle\ .$ (36)
Finally, we can re-interpret the degrees of freedom in these equations to live in a Hilbert space of dimension $(L+2)2^{n}$ with a clock register and an $n$-qubit working register, in which case the initial condition $\big\langle\dot{q}_{j}^{(l)}(0)\big\rangle=\delta_{l,1}\delta_{j,0}$ can be thought of as a quantum state $\left|l=1\right\rangle\otimes\left|0_{P}0_{Q}0_{R}\right\rangle$, which evolves (non-unitarily) under the equations (36).
B.3 Separation of Quantum Circuit Logic and Dynamics on the Clock Register
To analyze the matrix $\tilde{A}$ which generates the dynamics, we separate the quantum circuit logic, encoded in the $U_{l}$, from the hopping Hamiltonian acting on the clock register $\{\left|l\right\rangle\}_{l=1}^{L+2}$. This can be achieved through the basis transformation
$S$ $=\sum_{l=1}^{L+2}\left|l\right\rangle\left\langle l\right|\otimes U_{L+1}\cdots U_{l}\ .$ (37)
The transformed initial state reads $S(\left|l=1\right\rangle\otimes\left|0_{P}0_{Q}0_{R}\right\rangle)=\left|l=1\right\rangle\otimes(U\left|0_{P}0_{Q}0_{R}\right\rangle)$, with $U=\prod_{l=1}^{L+1}U_{l}$. It thus directly encodes the output state $U\left|0_{P}0_{Q}0_{R}\right\rangle$ of the computation. Note that
$U\left|0_{P}0_{Q}0_{R}\right\rangle=:\sqrt{1-\delta^{2}}\left|0_{P}\Psi_{0}\right\rangle+\delta\left|1_{P}\Psi_{1}\right\rangle\ ,$ (38)
where $\left|\Psi_{1}\right\rangle:=\tfrac{1}{\delta}\langle 1_{P}\rvert U\left|0_{P}0_{Q}0_{R}\right\rangle$ is the normalized postselected state which encodes the final result $\|\langle 0_{Q}|\Psi_{1}\rangle\|^{2}$ of the postselected circuit problem, with postselection probability $\delta^{2}=\Omega(2^{-n})$.
In the new basis, the matrix $\tilde{A}$ generating the dynamics becomes $X=S\tilde{A}S^{\dagger}$, where $X$ is of the form
$X=X_{0}\otimes\left|0_{P}\right\rangle\left\langle 0_{P}\right|\otimes\mathds{1}_{n-1}+X_{1}\otimes\left|1_{P}\right\rangle\left\langle 1_{P}\right|\otimes\mathds{1}_{n-1}\ ,\mbox{\ with\ }X_{i}=\begin{pmatrix}4&-1&\dots&0&0&0\\ -1&4&\dots&0&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ 0&0&\dots&4&-1&0\\ 0&0&\dots&-1&4&-1\\ 0&0&\dots&0&-1&4-\beta^{2}\delta_{i,1}\end{pmatrix}$. (39)
Here, $\mathds{1}_{n-1}$ denotes the identity operator on the remaining n1n-1 qubits (registers $Q$ and $R$). Importantly, we find that the entire dependence on the circuit $U$ has been absorbed into the basis change $S$, and the correspondingly updated initial state (39).
In the new basis, $X$ describes a dynamics which acts purely on the clock register, and which is controlled solely by the value of the postselected qubit $\{\left|0_{P}\right\rangle,\left|1_{P}\right\rangle\}$. The analysis of the dynamics thus reduces entirely to studying the properties of the clock operators $X_{0}$ and $X_{1}$. For the sake of the proof, we can absorb the transformation $S$ in a redefinition of coordinates $\dot{q}\to\dot{q}^{\prime}=S\dot{q}$, and we will henceforth work in this transformed basis. To understand the dynamics of the system, $\dot{q}^{\prime}(t)=\cos(\sqrt{X}t)\dot{q}^{\prime}(0)$, in the new basis, we now only need to analyze the properties of $X$.
Importantly, note that due to the way in which $S$ is constructed, the state of the computational registers $PQR$ at time $l=L+1$, which is what we choose for the read-out, is the same in the original and the transformed basis, such that we can analyze the read-out in the new basis.
B.4 Dynamics of the clock register: Tight binding model
The matrices $X_{0}$ and $X_{1}$ are finite Jacobi matrices. Intuitively, $X_{0}$ describes a single particle hopping on a one-dimensional chain; this can also be considered as the lattice discretization of a free particle on a finite line. $X_{1}$ differs from $X_{0}$ in that it additionally has a single-site potential well at the right end of the chain. If this well is deep enough (i.e., $\beta$ is large enough), it will thus possess an exponentially localized bound state at the right end of the chain. (Due to the lattice discretization, where the well is localized at a single site, there can only be one such bound state.)
More precisely, the properties of $X_{0}$ and $X_{1}$ and the implications on the resulting dynamics are as follows (see Section B.6 for the detailed derivation): The spectrum of $X_{0}$ lies in the interval $[2;6]$; the resulting time evolution operator $\cos(\sqrt{X_{0}}t)$, cf. Eq. (36), is therefore unitary. $X_{1}$ differs from $X_{0}$ precisely in the rank-one boundary perturbation $\beta^{2}\delta_{i,1}$ in the bottom right corner. It is well known that rank-one boundary perturbations to lattice Jacobians can produce at most one eigenvalue outside the spectral band of the unperturbed operator, and thus at most one bound state (i.e., a negative eigenvalue), see for example [16] or [17].
The scenario with such a negative eigenvalue is precisely the case we are interested in: If $X_{1}$ has such a negative eigenvalue, this will give rise to an exponentially growing eigenstate projector in the time evolution operator $\cos(\sqrt{X_{1}}t)=\cosh(\sqrt{-X_{1}}t)$, which will therefore dominate the full evolution $\cos(\sqrt{X}t)=\cos(\sqrt{X_{0}}t)\oplus\cos(\sqrt{X_{1}}t)$ at sufficiently long times, and thus amplify the postselected sector. As we show in Subsection B.6, the corresponding eigenvector is indeed exponentially localized at the right boundary, yet still has an at least exponentially small weight also at $l=1$. As we analyze in the next subsection, this ensures that the initial state (37), for which $\delta^{2}=\Omega(2^{-n})$, has sufficient overlap with the corresponding eigenvector to make the postselected sector dominant after propagation by a time linear in $L$, and to have sufficient weight on the time $l=L+1$.
While the Jacobi matrices $X_{0}$ and $X_{1}$ describe standard tight-binding models, which have well-known solutions, we show these required properties explicitly in Section B.6 for the sake of completeness.
B.5 Time Evolution and Read-Out
After evolution for a time $t$, the state $\big\langle\dot{q}(t)\big\rangle=\cos(\sqrt{X}\,t)\big\langle\dot{q}(0)\big\rangle$, expressed in the quantum register notation, will be
$\left|\Phi(t)\right\rangle$ $=\cos(\sqrt{X}\,t)\,\Big[\left|l=1\right\rangle\otimes U\left|0_{P}0_{Q}0_{R}\right\rangle\Big]$ (40)
[…]
$=\sqrt{1-\delta^{2}}\,\Big[\cos(\sqrt{X_{0}}\,t)\,\left|l=1\right\rangle\Big]\otimes\left|0_{P}\Psi_{0}\right\rangle+\delta\,\Big[\cos(\sqrt{X_{1}}\,t)\,\left|l=1\right\rangle\Big]\otimes\left|1_{P}\Psi_{1}\right\rangle\ .$ (42)
We now use the following: First, $X_{0}$ is positive semi-definite, and therefore $\cos(\sqrt{X_{0}}\,t)$ is unitary and thus norm-preserving. Second, we choose $\beta$ such that $X_{1}$ has precisely one negative eigenvalue $\alpha_{1}<0$, with corresponding eigenvector $\left|\chi_{1}\right\rangle$, see Lemma 1 in Section B.6; thus, $\cos(\sqrt{X_{1}}\,t)=\cosh(\sqrt{-\alpha_{1}}\,t)+V(t)$, where $V(t)$ is once again unitary, i.e. norm-preserving. Together, this yields the estimate
$\left|\Phi(t)\right\rangle$ $=\delta\,\cosh(\sqrt{-\alpha_{1}}\,t)\,\left|\chi_{1}\right\rangle\langle\chi_{1}|l=1\rangle\otimes\left|1_{P}\Psi_{1}\right\rangle+O(1)\ .$ (43)
Let us now define
$\left|\Phi^{\prime}(t)\right\rangle:=$ $\>\big(\left\langle l=L+1\right|\otimes\left\langle 1_{P}\right|\big)\,\left|\Phi(t)\right\rangle$ (44)
$=$ $\>\underbrace{\delta\,\cosh(\sqrt{-\alpha_{1}}\,t)\,\langle l=L+1\left|\chi_{1}\right\rangle\langle\chi_{1}|l=1\rangle}_{=:W(t)}\otimes\left|\Psi_{1}\right\rangle+O(1)\ ,$ (45)
the state obtained after projecting $\left|\Phi(t)\right\rangle$ onto time step $l=L+1$ and the state $\left|1_{P}\right\rangle$ to be postselected on. In this notation, the read-out (32) then amounts to measuring $\zeta(I_{\mathrm{den}};t)=\big\|\left|\Phi^{\prime}(t)\right\rangle\big\|^{2}$ and $\zeta(I_{\mathrm{num}};t)=\big\|\langle 1_{Q}\left|\Phi^{\prime}(t)\right\rangle\big\|^{2}$, respectively. We required that our measurements gave estimates with $O(1)$ accuracy; this finite precision can be readily absorbed in the $O(1)$ term in (45) and thus does not need to be accounted for separately.
Before putting things together, we still need to bound the prefactor $W(t)$ introduced in Eq. (45) from below. In Lemma 1 in Section B.6, we show that for $\beta>2$, we have $\alpha_{1}=2\cosh(\kappa_{1})-4\geq\beta^{2}-4>0$, $|\langle l=1|\chi_{1}\rangle|^{2}=\Omega\big(\tfrac{1}{L+2}e^{-2L\kappa_{1}}\big)$, and $|\langle l=L+1|\chi_{1}\rangle|^{2}=\Omega(1)$. Combining this with $\delta^{2}=\Omega(2^{-L})$ and $\cosh(x)^{2}=\Omega(e^{2x})$, we obtain
$|W(t)|^{2}=\Omega\Big(\tfrac{1}{L+2}\,e^{2\sqrt{-\alpha_{1}}\,t-(2\kappa_{1}+\log 2)\,L}\Big)$ (46)
and thus, for any $\beta>2$, there are constants $c>0$, $c^{\prime}>0$ such that for $t_{f}\geq cL=O(L)$, it holds that $W(t_{f})=\Omega(e^{c^{\prime}L})$.
We are now ready to analyze the read-out. The algorithm accepts if
$\frac{1}{2}<\frac{\zeta(I_{\mathrm{num}};t_{f})}{\zeta(I_{\mathrm{den}};t_{f})}=\frac{\big\|W(t_{f})\langle 1_{Q}|\Psi_{1}\rangle+O(1)\big\|^{2}}{\big\|W(t_{f})\,\lvert\Psi_{1}\rangle+O(1)\big\|^{2}}=\|\langle 1_{Q}|\Psi_{1}\rangle\|^{2}+O\big(\tfrac{1}{W(t_{f})}\big)=\|\langle 1_{Q}|\Psi_{1}\rangle\|^{2}+O(e^{-cL})$ (47)
(where we have used $\|\,\lvert\Psi_{1}\rangle\|=1$), that is, it accepts exactly if $\|\langle 1_{Q}|\Psi_{1}\rangle\|^{2}>2/3$ (vs. <1/3<1/3), i.e. exactly for accepting instances of the postselected circuit problem. This completes the proof.
B.6 Solution of the Tight-Binding Model on the Clock Register
The purpose of this Section is to establish the technical bound underlying the amplitude separation induced by the final-layer perturbation. We solve the spectrum of the matrices $X_{0}$ and $X_{1}$, defined in Section B.3. The spectral properties of the matrices $X_{0}$ and $X_{1}$ follow from standard results for finite Jacobi matrices with boundary perturbations. We include the derivation here for completeness and to make all constants explicit.
The two blocks $X_{0}$ and $X_{1}$, which act on the subspace of the post-selected qubit in the state $\left|0_{P}\right\rangle$ or $\left|1_{P}\right\rangle$ respectively, can be solved separately. The eigenvectors and -values of $X_{0}$ are the same as in [2], i.e.
$X_{0}\left|\phi_{l}\right\rangle=\gamma_{l}\left|\phi_{l}\right\rangle,$
$\text{with}\quad\left|\phi_{l}\right\rangle=\sqrt{\frac{2}{L+3}}\sum_{l^{\prime}=1}^{L+2}\sin\left(\frac{ll^{\prime}\pi}{L+3}\right)\left|l^{\prime}\right\rangle\quad\text{and}\quad\gamma_{l}=4-2\cos\left(\frac{\pi l}{L+3}\right).$ (48)
The operator $X_{1}$ acts on the subspace of $\left|1_{P}\right\rangle$ similar to $X_{0}$, but with an additional term $\beta^{2}\left|L+2\right\rangle\left\langle L+2\right|$. The spectrum of $X_{1}$ can be solved analogously to the $\beta=0$ case using standard methods. The eigenvectors $\left|\psi_{l}\right\rangle=\sum_{l^{\prime}=1}^{L+2}\psi_{ll^{\prime}}\left|l^{\prime}\right\rangle$ satisfy the difference equations
$4\psi_{ll^{\prime}}-\psi_{l(l^{\prime}+1)}-\psi_{l(l^{\prime}-1)}$ $=\alpha_{l}\psi_{ll^{\prime}}\quad\forall l^{\prime}\in\{1,...,L+1\}$ (49)
$(4-\beta^{2})\psi_{l(L+2)}-\psi_{l(L+1)}$ $=\alpha_{l}\psi_{l(L+2)}$ (50)
using $\psi_{l0}=0$ for compact notation. The difference equations are solved by the following ansatz $\psi_{ll^{\prime}}=c_{l}\sin(l^{\prime}\omega_{l})$ and $\alpha_{l}=4-2\cos(\omega_{l})$. This can be quickly checked by inserting into Eq. (49) (omitting the normalization constant $C$)
$4\sin(l^{\prime}\omega_{l})-\sin((l^{\prime}+1)\omega_{l})-\sin((l^{\prime}-1)\omega_{l})$
$=4\sin(l^{\prime}\omega_{l})-\sin(l^{\prime}\omega_{l})\cos(\omega_{l})-\cos(l^{\prime}\omega_{l})\sin(\omega_{l})-\sin(l^{\prime}\omega_{l})\cos(\omega_{l})+\cos(l^{\prime}\omega_{l})\sin(\omega_{l})$
$=(4-2\cos(\omega_{l}))\sin(l^{\prime}\omega_{l}),$ (51)
using the angle addition theorem, $\sin(x\pm y)=\sin x\cos y\pm\cos x\sin y$ for all $x,y\in\mathds{C}$. Inserting the ansatz into Eq. (50) yields a quantization condition on the frequencies $\omega_{l}$
$(2\cos(\omega_{l})-\beta^{2})\sin((L+2)\omega_{l})$ $=\sin((L+2-1)\omega_{l})=\sin((L+2)\omega_{l})\cos(\omega_{l})-\cos((L+2)\omega_{l})\sin(\omega_{l})$
$\iff\beta^{2}=\cos(\omega_{l})+\frac{\sin(\omega_{l})}{\tan((L+2)\omega_{l})}.$ (52)
In the $\beta=0$ case, this yields $\tan(\omega_{l})=\tan((L+2)\omega_{l})$, which determines the frequencies used above. For nonzero $\beta$, there are no analytical solutions to this equation, in general. However, we will show that, for large enough $\beta$, there exists a unique (pair of) solution(s) of $\omega_{1}=\pm i\kappa_{1}$ to Eq. (52) that is purely imaginary and ensures a negative eigenvalue $\alpha_{1}=4-2\cosh(\kappa_{1})<0$, and ultimately creates exponential growth in $\cos(\sqrt{X_{1}}t)$ as opposed to oscillatory dynamics from positive $\alpha_{l},\gamma_{l}$.
Lemma 1.
Let $\kappa_{1}=i\omega_{1}\in\mathds{R}$ be a solution of Eq. (52) for $\beta>2$ and $L\geq 0$.
  1. 1.
    There exists a unique negative eigenvalue $\alpha_{1}=4-2\cosh(\kappa_{1})$, separated from zero by a constant, $\alpha_{1}\leq 4-\beta^{2}<0$.
  2. 2.
    The corresponding eigenvector $\left|\psi_{1}\right\rangle$ has an overlap $\left\lvert\left\langle l=1|\psi_{1}\right\rangle\right\rvert^{2}=\Omega\left(\frac{e^{-2L\kappa_{1}}}{L+2}\right)$.
  3. 3.
    The corresponding eigenvector $\left|\psi_{1}\right\rangle$ has an overlap $\left\lvert\left\langle l=L+1|\psi_{1}\right\rangle\right\rvert^{2}=\Omega(1)$ that is separated from zero by a constant.
Proof.
  1. 1.
    We start from Eq. (52) and bound
    $\cosh(\kappa_{1})=\beta^{2}-\sinh(\kappa_{1})\coth((L+2)\kappa_{1})=\beta^{2}-\tanh(\kappa_{1})\cosh(\kappa_{1})\coth((L+2)\kappa_{1}),$ (53)
    substituting $\sinh(\kappa_{1})=\tanh(\kappa_{1})\cosh(\kappa_{1})$. Now, we solve again for $\cosh(\kappa_{1})$ to get
    $\cosh(\kappa_{1})=\frac{\beta^{2}}{1+\tanh(\kappa_{1})\coth((L+2)\kappa_{1})}\geq\frac{\beta^{2}}{2},$ (54)
    where we used $\tanh(x)\coth((L+2)x)\leq 1$ for $L\geq 0$. Uniqueness of $\alpha_{1}$ is a consequence of the reality of the eigenvalues of $X_{1}$; $\alpha=4-2\cos(\omega)$ is real if, and only if $\omega$ is purely real or purely imaginary, otherwise $\cos(x+iy)=\cos(x)\cosh(y)-i\sin(x)\sinh(y)$ is not real. For $\omega$ real, $\alpha$ cannot be negative, hence the only possible solution is $\omega=\pm i\kappa$, for which there is only one solution for $\kappa$ through Eq. (52) given a value for $\beta^{2}$, which is a monotonic function in $\kappa$. Hence exactly one exponentially growing mode exists in the $\left|1_{P}\right\rangle$ sector, while all other modes remain oscillatory.
  2. 2.
    Consider the normalized eigenvector corresponding to the eigenvalue $\alpha_{1}$
    $\left|\psi_{1}\right\rangle=c_{1}\sum_{l^{\prime}=1}^{L+2}\sinh(l^{\prime}\kappa_{1})\left|l^{\prime}\right\rangle\qquad\text{with}\qquad c_{1}=\left(\sum_{l^{\prime}=1}^{L+2}\sinh(l^{\prime}\kappa_{1})^{2}\right)^{-\frac{1}{2}}=2\left(\frac{\sinh((2L+5)\kappa_{1})}{\sinh(\kappa_{1})}-(2L+5)\right)^{-\frac{1}{2}},$ (55)
    We bound the overlap with the $l=1$ using the inequality $c_{1}\geq\frac{1}{\sqrt{L+2}}\cdot\frac{1}{\sinh((L+2)\kappa_{1})}$. Thus,
    $\left\lvert\left\langle l=1|\psi_{1}\right\rangle\right\rvert$ $\geq\frac{1}{\sqrt{L+2}}\frac{\sinh(\kappa_{1})}{\sinh((L+2)\kappa_{1})}=\Theta\left(\frac{e^{-L\kappa_{1}}}{\sqrt{L+2}}\right).$ (56)
  3. 3.
    The third statement can be similarly bounded. Using partial sums of the geometric series, one can calculate $c_{1}=2\left(\frac{\sinh((2L+5)\kappa_{1})}{\sinh(\kappa_{1})}-(2L+5)\right)^{-\frac{1}{2}}$. With this, we can bound
    $\left\lvert\left\langle l=L+1|\psi_{1}\right\rangle\right\rvert^{2}=2\frac{\sinh((L+1)\kappa_{1})^{2}}{\frac{\sinh((2L+5)\kappa_{1})}{\sinh(\kappa_{1})}-(2L+5)}\geq 2\sinh(\kappa_{1})\frac{e^{2(L+1)\kappa_{1}}-2}{e^{(2L+5)\kappa_{1}}}\geq 2\sinh(\kappa_{1})e^{-3\kappa_{1}}(1-2e^{-2\kappa_{1}}),$ (57)
    by dropping positive terms in the numerator and negative terms in the denominator. This lower bound has a constant separation from zero.
The crude bound $\tanh(x)\coth((L+2)x)\leq 1$ in the proof of statement one of Lemma 1 leads us to the constraint $\beta>2$, but can be tightened to the $L$-dependent bound
$\beta^{2}>2+\sqrt{3}\;\frac{(2+\sqrt{3})^{2(L+2)}+1}{(2+\sqrt{3})^{2(L+2)}-1},$ (58)
using the critical value $\kappa_{0}=\cosh^{-1}(2)=\log(2+\sqrt{3})$. Eq. (58) includes the value $\beta=2$, for instance, to ensure a constant separation of $\alpha_{1}$ from zero. Using $\kappa_{1}\geq\kappa_{0}$, also yields a numerical lower bound of $5\%$ for the third statement of Lemma 1.
B.7 Alternative PostBQP-hard families of bosonic Hamiltonians
At the end of this appendix, let us argue that the hardness result is robust under more general choices of $E^{+}$. One might wonder how restrictive the choice of nonzero $E^{+}$ is for $\mathsf{PostBQP}$-hardness. Indeed, $\mathsf{PostBQP}$-hard instances can be found in a more general class of Hamiltonians. Off-diagonal terms $q_{j}p_{k}+q_{k}p_{j}$ of $E^{+}$ can be diagonalized by changing the qubit encoding using a symplectomorphism $q\mapsto Oq$ and $p\mapsto(O^{-1})^{T}p$. If we were to only switch on a single term, post-selection would not target a single-qubit output $\left|1_{P}\right\rangle$ but rather a whole bitstring, which is equivalent to post-selecting only one qubit at the end and can be mapped onto another using an ancillary register [15].
Another family of Hamiltonians whose simulation is $\mathsf{PostBQP}$-hard is given by the constraint $E^{+}=0$. As we show in appendix C, besides $E^{+}$ also $\Delta=A-C$ can generate non-unitary dynamics. Setting $E=iE^{-}$, the dynamics is governed by the differential equation
$\ddot{q}=-iE^{-}\dot{q}+C\dot{p}=-\left((E^{-})^{2}+CA\right)q-i(E^{-}C+CE^{-})p.$ (59)
If we can get rid of the term proportional to $p$, we can reproduce Eq. (33), in particular the differential operator $\tilde{A}$. In particular, we have to choose $A,C$ real symmetric and $E^{-}$ real antisymmetric, such that
  1. 1.
    $E^{-}C+CE^{-}=0$
  2. 2.
    CACA acts as a Feynman-Kitaev Hamiltonian
  3. 3.
    $(E^{-})^{2}=-\beta^{2}\Pi$ with an adequate projector $\Pi$ onto the postselected part of the readout layer
These conditions can be implemented, for instance, by defining two Feynman-Kitaev dynamics on a doubled system
$A$ $=A_{FK}\otimes Z$ (60)
$C$ $=\mathds{1}\otimes Z$ (61)
$E^{-}$ $=\beta\left|L+2\right\rangle\left\langle L+2\right|\otimes\left|1_{P}\right\rangle\left\langle 1_{P}\right|\otimes(iY)$ (62)
with $A_{FK}$ being the matrix defined in Eq. (28) and $Z$ and $Y$ are Pauli matrices. One can directly check that (1.) $E^{-}$ and $C$ anticommute, (2.) $CA$ describes two copies of the same Feynman-Kitaev Hamiltonian and (3.) $(E^{-})^{2}=-\beta^{2}\Pi_{L+2,1_{P}}\otimes\mathds{1}$ postselects onto the last readout layer in both copies.
$(E^{-})^{2}+CA=\tilde{A}\otimes\mathds{1},$ (63)
reproduces the generator used in the proof above. It has the same spectral properties – as before – and when starting in the analog initial state with only $\left\langle p^{(l=1)}_{0}(0)\right\rangle=1$, one can solve the postselected circuit problem by estimation of the momenta in the output layer. Note that while $A$ and $C$ individually are indefinite, the product $CA$ only has positive eigenvalues, which shows that the non-unitary dynamics necessary for postselection comes solely from $(E^{-})^{2}$, cf. also appendix C.2.
A third family of Hamiltonians which are $\mathsf{PostBQP}$-hard is given by $E=0$, but lifting the condition that $CA$ has to have a positive spectrum, as we require in systems of inertially coupled bosons. In this family of Hamiltonians one can directly choose $A$ to be the targeted generator $\tilde{A}$ as in Eq. (33) with a negative eigenvalue, which is now allowed.
Appendix C Coordinate Transformations
This appendix analyzes the conditions under which the first-moment dynamics of quadratic bosonic systems remain bounded, i.e., do not exhibit exponential amplification. While the main text focuses on inertially coupled bosons, the following discussion treats general quadratic Hamiltonians and later specializes to the families shown in Fig. 1. Finally, we demonstrate how coordinate transformations can be used to interchange terms in the resulting differential equations, thereby identifying the boundaries between simulable families of Hamiltonians.
C.1 Most General Quadratic Bosonic Hamiltonian
The most general quadratic Hamiltonian can be written as
$H=\frac{1}{2}\begin{pmatrix}q&p\end{pmatrix}\begin{pmatrix}A&E\\ E^{T}&C\end{pmatrix}\begin{pmatrix}q\\ p\end{pmatrix}$, (64)
where the matrix $E=E^{+}+iE^{-}$ can be split into a real and symmetric part $E^{+}$ contributing with terms $\frac{1}{2}E_{jk}(q_{j}p_{k}+q_{k}p_{j})$ and an imaginary, antisymmetric (i.e. Hermitian) part $E^{-}$ contributing with terms of the form $\frac{1}{2}E_{jk}(q_{j}p_{k}-q_{k}p_{j})$. In matrix form, the equations of motion become
$\partial_{t}\begin{pmatrix}iq\\ p\end{pmatrix}=-i\begin{pmatrix}iE^{T}&-C\\ -A&-iE\end{pmatrix}\begin{pmatrix}iq\\ p\end{pmatrix}=-i\left(\mathds{1}\otimes E^{-}+(iZ)\otimes E^{+}+(iY)\otimes\Delta-X\otimes\Sigma\right)\begin{pmatrix}iq\\ p\end{pmatrix}$, (65)
with $\Sigma=\frac{1}{2}(A+C)$ and $\Delta=\frac{1}{2}(A-C)$. For the first-moment dynamics to remain bounded—i.e., to avoid exponential amplification or suppression of the variables of interest—the matrix in Eq. (65) must have purely real eigenvalues, corresponding to oscillatory eigenmodes. A sufficient condition ensuring this is that the generator in Eq. (65) is anti-Hermitian.
C.2 Positivity of coupling matrices $A$ and $C$
As we show in the main text, coordinates can be chosen as $B^{\dagger}q$ and $D^{\dagger}p$ for $A=BB^{\dagger}$ and $C=DD^{\dagger}$, such that correlations evolve unitarily in the absence of mixed terms, $E=0$. If in addition to positivity of $A$ and $C$, the matrix elements are derived from positive coupling constants (cf. Eqs. (17) and (18)), an explicit and efficiently queriable form of $B$ and $D$ is known, cf. Eq. (7). If this sign convention is not met, an oracle access to quantum circuits implementing $\sqrt{A}$ and $\sqrt{C}$ can be employed to simulate time evolution via quantum phase estimation, cf. Theorem 4 of [2].
In fact, the condition that both, $A$ and $C$, are PSD is stronger than necessary. In the case $E=0$, a positive spectrum of the product $AC$ is sufficient to ensure unitary evolution, cf. Eq. (11). This can be achieved using indefinite or even negative $A$ and $C$ if they share the same block structure in a decomposition into positive and negative part such that signs are correctly canceled in the product. An example that deviates from this condition is the single-mode squeezing Hamiltonian
$H=\frac{1}{2}(q^{2}-p^{2})=\frac{1}{2}(a^{2}+(a^{\dagger})^{2}),$ (66)
corresponding to $A=1$ and C=1C=-1. In this case the dynamics exponentially amplifies one quadrature, while suppressing the conjugate one. Effectively, this projects our encoded state $\left|\psi\right\rangle$ onto the squeezed quadrature eigenstate, $\frac{1}{\sqrt{2}}\left(\left|\alpha=1\right\rangle+\left|\alpha=\bar{1}\right\rangle\right)$.
C.3 Symplectic Transformations and Generalized Hopping
Upon inspection of Eq. (65), $R$-point correlations of $\left\langle iq\right\rangle$ and $\left\langle p\right\rangle$ can also be directly simulated (without transforming into the coordinates $B^{\dagger}q$ and $D^{\dagger}p$) by switching off the non-Hermitian terms $E^{+}=\Delta=0$. This extends the family of quantum walk Hamiltonians discussed in the main text (cf. Eq. (16)) by extending the graph to complex weights, where imaginary parts correspond to $E^{-}\neq 0$.
Using unitary transformations on the Paulis in Eq. (65) allows for a permutation of terms and permutations of the phase factor $i$, which determines Hermiticity. In particular, a symplectic transformation, that is a coordinate transformation which leaves the symplectic form $\Omega=(iY)\otimes\mathds{1}$ unchanged, also leaves the equations of motion invariant. One example is the transformation
$-i\exp\left(-i\frac{\pi}{4}Y\right)\otimes\mathds{1}=\frac{-i}{\sqrt{2}}\begin{pmatrix}\mathds{1}&-\mathds{1}\\ \mathds{1}&\mathds{1}\end{pmatrix}$, (67)
which maps the $(iq,p)$ basis onto the Fock basis spanned by ladder operators $(a,a^{\dagger})$. These coordinates were used in the main text (cf. discussion around Eq. (15)) to show the equivalence of this simulable family, $E^{+}=\Delta=0$ to number-preserving Hamiltonians.